OpenStudy (anonymous):

Simplify the expression (3n^2)^3

OpenStudy (anonymous):

ok, do you have any clue on how to solve this?

OpenStudy (anonymous):

Yeah I just don't know if 3 is raised to the 3rd power

OpenStudy (anonymous):

\[(3n ^{2})^{3}\] = \[3^{3}n ^{2 \times 3} \] = \[27n ^{6}\]

OpenStudy (anonymous):

yep, it is.

OpenStudy (anonymous):

do you understand my working?

OpenStudy (anonymous):

Yea, can you help me with this one (3m^1/2•27n^1/4)^4

OpenStudy (anonymous):

it's that, correct?

OpenStudy (anonymous):

No it's 3m^1/2 and 27n^1/4

OpenStudy (anonymous):

Yea sorry I'm just on my phone so it's a bit slow, it's like this (3m ^(1/2) •27n^(1/4)^4

OpenStudy (anonymous):

\[3m ^{1/2} \times 27n ^{1/4} \] \[3m ^{1/2} \times 27n ^{1/256} \] \[\sqrt{3m}.\sqrt[256]{27n}\]

OpenStudy (anonymous):

The first step I was unable to type ^4 additionally to the (1/4) power, sorry, but as you can see I have expanded as normal with the ^4.

OpenStudy (anonymous):

@robtobey, is my working correct?

OpenStudy (anonymous):

I got\[81 \sqrt{m}* n^{1/256} \]

OpenStudy (anonymous):

yeh, u should be right.

OpenStudy (anonymous):

How did you get 1/256?

OpenStudy (anonymous):

(1/4)^4

OpenStudy (anonymous):

\[(3*m){}^{\wedge}(1/2)*27*n{}^{\wedge}(1/4){}^{\wedge}4=27 \sqrt{3} *\sqrt{m}* n^{1/256} \]

OpenStudy (anonymous):

\[3m ^{1/2} \times 27n ^{1/4} \] \[m ^{1/2} = \sqrt{m}\] \[81\sqrt{m}*n ^{1/256}\]

OpenStudy (anonymous):

Oh okay I see now thank you